Car A of mass of 2000 kg moving at 72 km/h. east try to overtake car B of mass


1250 kg moving at 60 km/h, east. Car A collides with car B. The car stick


together after the collision. Find their velocity after collision. (It should be in


m/s).

1
Expert's answer
2021-11-07T19:26:31-0500

"m_1v_1+m_2v_2=u(m_1+m_2)," "\implies"

"u=\frac{m_1v_1+m_2v_2}{m_1+m_2}=67.4~\frac{km}h=18.7~\frac ms."